OpenStudy (anonymous):

A truck's odometer reads 69,189.7 at the beginning of a trip and 69,476.7 at the end. If the trip takes 14.1 gallons of gas, what is the gas mileage in miles per gallon? (Round to the nearest tenth)

OpenStudy (anikhalder):

No of miles travelled= 69476.7-69189.7=287 miles Gas consumed=14.1 gallons Hence, On 14.1 gallons a trip of 287 miles is made On 1 gallon a trip of 287/14.1= 20.35 miles/gallon To the nearest tenth, the gas mileage is 20.4 miles/gallon